OpenStudy (anonymous):

rewrite the expression whithout using grouping symbols 7(x-19)

Vocaloid (vocaloid):

hint: use the distributive property know where to start?

OpenStudy (anonymous):

lol i dont realy no how

Vocaloid (vocaloid):

hint: A(B-C) = A*B - A*C follow the pattern 7(x-19) = ?

OpenStudy (anonymous):

is it 7x-7 times 19

Vocaloid (vocaloid):

nope not quite

Vocaloid (vocaloid):



Vocaloid (vocaloid):



OpenStudy (anonymous):

7x minus 19

Vocaloid (vocaloid):

again, not quite...

Vocaloid (vocaloid):

A matches up with 7, and B matches up with x, and C matches up with 19 so what is A*B - A*C?

OpenStudy (anonymous):

7x minus 19

Vocaloid (vocaloid):

nope...

Vocaloid (vocaloid):

A = 7 B = x C = 19 what is A*B - A*C?

OpenStudy (anonymous):

12

Vocaloid (vocaloid):

nope...

Vocaloid (vocaloid):

A = 7 B = x what is A*B?

OpenStudy (anonymous):

i dont no lol

Vocaloid (vocaloid):

think carefully... A is 7 B is x what is A times B?

Vocaloid (vocaloid):

just multiply A and B together.......

OpenStudy (anonymous):

what does x equal

Vocaloid (vocaloid):

x is an unknown

Vocaloid (vocaloid):

since A is 7, and B is x, A times B is 7x, with me so far?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

26

Vocaloid (vocaloid):

let me type that again

Vocaloid (vocaloid):

A = 7 C = 19 what is A*C?

OpenStudy (anonymous):

what does the star stand for

Vocaloid (vocaloid):

* means multiply

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

133

Vocaloid (vocaloid):

good. so if A*B = 7x and A*C = 133 then what is A*B - A*C?

OpenStudy (anonymous):

126